Let A be a 3 x 3 matrix A with eigenvalues
1, –2, 2. Find the trace of A + A². Give
reasons to justify your answer.
1
Expert's answer
2020-05-11T16:37:29-0400

Eigen values of A are 1,-2,2 (Given).

Therefore eigen values of A2 are 1, 4, 4.

So eigen values of (A+A2) are 2, 2, 6.

Trace of (A+A2)=Sum of eigen values of (A+A2)

= 2+2+6=10.

  • Answer is 10.